lululara:

Kate is making pizza. She puts 8 ounces of cheese on each pizza. If she has 4 24-ounce packages of cheese, how many pizzas can she make?

umm:

Welcome to QuestionCove. If she puts 8 ounces of cheese on each pizza's and she has 24 ounce packages, how many times do you think 8 can go into 24?

jhonyy9:

8 ounces of cheese on each pizza's she has 4 of 24-ounce packages cheese 4*24 = 96 ounces total cheese 8 ounces 1 pizza's 96 ounces x pizzas's x = 96*1/8 = 96/8 = 12 answer : she can make 12 pizzas hope helped